A bicyclist rides 5.0 km due east, while the resistive forcefrom the air has a magnitude of 3.0 N and points due west. Therider then turns around and rides 5.0 km due west, back to herstarting point. The resistive force from the air on the return triphas a magnitude of 3.0 N and points due east.
a) find the work done by the resistive force during the roundtrip.
Based on answer in part A.
b)is the resistive force a conservative force? explain.
Answer #1
(a) W = -Fs-Fs = -2(3N)(5 X 10^3 m) = -3 X 10^4 J
(b) Not conservative because it does net work on thebicyclist